# Advances in 4D Representation: Geometry, Motion, and Interaction

### v1

@@ Abstract
We present a survey on 4D generation and reconstruction, a fast-evolving subfield of computer graphics whose developments have been propelled by recent advances in neural fields, geometric and motion deep learning, as well 3D generative artificial intelligence (GenAI). While our survey is not the first of its kind, we build our coverage of the domain from a unique and distinctive perspective of 4D representations, to model 3D geometry evolving over time while exhibiting motion and interaction. Specifically, instead of offering an exhaustive enumeration of many works, we take a more selective approach by focusing on representative works to highlight both the desirable properties and ensuing challenges of each representation under different computation, application, and data scenarios. The main take-away message we aim to convey to the readers is on how to select and then customize the appropriate 4D representations for their tasks. Organizationally, we separate the 4D representations based on three key pillars: geometry, motion, and interaction. Our discourse will not only encompass the most popular representations of today, such as neural radiance fields (NeRFs) and 3D Gaussian Splatting (3DGS), but also bring attention to relatively under-explored representations in the 4D context, such as structured models and long-range motions. Throughout our survey, we will reprise the role of large language models (LLMs) and video foundational models (VFMs) in a variety of 4D applications, while steering our discussion towards their current limitations and how they can be addressed. We also provide a dedicated coverage on what 4D datasets are currently available, as well as what is lacking, in driving the subfield forward.
@@ 1 Introduction
Recent surveys in this domain have primarily categorized methods by applications, approaches, and/or the granularity of the extracted scene signals. Cao et al. [CLH25] provides a structured review of recent progress in reconstructing 4D spatial intelligence, which is defined as understanding 3D scenes and their evolution over time, advancing from extracting low-level geometric cues such as depth, pose, and point maps toward complex reasoning encompassing interactions and physics. This survey does highlight the growing importance of advanced 3D representations while focusing mainly on neural radiance fields (NeRFs) and 3D Gaussian splatting (3DGS). Along the same lines, Fan et al. [FZZ25] presents a comprehensive survey of recent progress in dynamic scene representation and reconstruction, focusing on the transition from NeRF to 3DGS and their trade-offs. This second survey systematically categorizes the works covered based on motion representation paradigms including rigid, articulated, non-rigid, and hybrid motions, and examines strategies for modeling temporal changes in geometry and appearance. In addition, Miao et al. [MLQ25] surveys the emerging field of 4D generation by introducing a taxonomy of low-level geometry representations (meshes, NeRFs, point clouds, and 3DGS), foundational techniques (diffusion models and score distillation sampling), and conditioning methods (text, images, videos, 3D inputs, and multimodal control). This survey further categorizes the algorithmic approaches (end-to-end, via generated data, implicit distillation, vs. explicit supervision) and highlights the growing range of applications for 4D generation.
@@
Organizationally, we separate the 4D representations based on three key pillars: geometry, motion, and interaction. We distinguish between structured representations, i.e., those that maintain explicit primitive or part delineations and relations, and unstructured representations, i.e., those that encode 4D content through implicit functions, point distributions, or learned feature spaces without explicit structural constraints. While the latter category encompasses the most popular representations of today, namely NeRFs and 3DGS, they are both built on rendering primitives, mainly for novel view synthesis. They are not best suited to prevalent 4D tasks such as modeling, editing, or interactions, for which more compact and structural representations, built on higher-level primitives, are more appropriate. Hence, we also discuss these representational choices even though they have been been relatively under-explored. Throughout our survey, we will reprise the role of large language models (LLMs) and video foundational models (VFMs) in a variety of 4D applications, while steering our discussion towards their current limitations and how they can be addressed. We also provide a dedicated coverage on what 4D datasets are currently available, as well as what is lacking, in driving the subfield forward.
@@
Our survey is organized as follows. Section˜2 presents our core representational analysis, examining unstructured representations (mesh, point clouds, NeRFs and 3DGS) and structured representations (template, primitives, and graphs), respectively. Section˜3 focuses on motion and temporal dynamics, analyzing how different motion types interact with representational choices. Section˜4 covers representation choices in modeling multiple entities that interact with each other. Section˜5 surveys datasets, evaluation metrics, and benchmarking frameworks that enable systematic comparison across representations. Finally, Section˜8 provides an overall analysis across representations and discusses emerging trends, persistent challenges, and future research directions in 4D representation. The overall taxonomy of our survey is shown in Figure˜2.
@@ 2 Modeling Geometry
We categorize representations into unstructured (Section˜2.1 and structured (Section˜2.2) classes on whether their operational primitives carry explicit functional, hierarchical, or semantic meaning. See Figure˜1 for an illustration. Unstructured representations—meshes (Section˜2.1.1), point clouds (Section˜2.1.2), NeRF (Section˜2.1.3), and Gaussian Splatting (Section˜2.1.4)—use primitives (vertices, points, ray samples, Gaussians) as independent geometric elements. While local connectivity may exist (e.g., mesh faces), these representations do not impose inter-primitive global structural constraints or functional decompositions.
@@ 2.1.1 Mesh-Based Representations
Mesh is mainly adopted in object-centric 4D modeling, target to produce animated mesh asset from text [DSW25, CHC24], video [CZTW25, LCL24] or static 3D asset [WYWB25, SLW25]. However, mesh representations face several challenges in 4D applications, which we discuss below alongside methods developed to address them.
Spatiotemporal Consistency: Animated mesh sequences require precise vertex-to-vertex correspondence across frames to preserve consistent texture and topology, preventing naive per-frame generation pipelines that can lead to different vertex-connectivity in each generated sample. V2M4 [CZTW25] circumvents this through a reconstruction-and-refine pipeline: it first applies image-to-3D reconstruction models [XLX25, ZLL25, LZL25, LLY24] to generate per-frame meshes, then sets the first frame as anchor and deforms it to register with subsequent frames, maintaining consistent mesh correspondence through continuous deformation.
@@
Open Challenges. Despite recent progress, fundamental limitations remain. Topological inflexibility prevents meshes from representing splitting or merging. Their inflexible topology and connectivity struggles with volumetric phenomena such as clouds, smoke, and fire, motivating the adoption of topologically flexible alternatives such as NeRF (Section˜2.1.3) and point clouds (Section˜2.1.2) for such scenarios. Additionally, current mesh-centric generation and reconstruction methods predominantly focus on object-level rather than scene-level modeling. The later has recently be studied on static 3D scene reconstruction [YZY25] through generative method, while most existing works relies on mesh database retrival [WIR24]. Key difficulties include: (1) generating and tracking separate mesh topologies for multiple objects at varying scales, (2) handling inter-mesh physical constraints, and (3) maintaining temporal coherence when objects move independently with different deformation patterns. Addressing these challenges remains an open problem for the research community.
@@ 2.1.2 Point-cloud-based Representation
A point cloud represents 3D geometry as an unordered set of points $P=\{p_{i}\}_{i=1}^{N}$, where each point $p_{i}$ is defined by spatial coordinates $(x,y,z)$ and optional attributes including color (RGB), intensity, surface normals, and timestamps. Their unstructured, unordered nature enables representation of arbitrary and changing topologies without connectivity constraints. Point clouds also serve as the primary output format for real-world acquisition mechanisms. This direct sensor correspondence has established point clouds as the standard representation for large-scale scene capture, particularly in autonomous driving (KITTI [GLSU13], nuScenes [CBL20]), where their geometric scalability outperform alternatives. On native 4D point cloud sequences, numerous works focus on improving point cloud quality through denoising [HH25], upsampling [LLF21, BBBB23], and mesh reconstruction [TADTBS22, RBZ20, NMOG19]. Another research direction addresses dynamic point cloud understanding, including motion recognition [LHL25, DZLY23], object detection and tracking [HGH22, WNGO24, CLZ24], scene flow estimation [LQG19, WWL20, TD20], forecasting [KHHR23, WYW24], and motion interpolation [ZQZ22, ZWL23, LCQ21]. With the emergence of large feed-forward geometric models like DUST3R [WLC24] and VGGT [WCK25], a new promising branch of works is to create dynamic point maps from videos [ZHH24, WZH25, WZZ25, TZW25] and incorporating tracking within the inferred point maps [XWX25].
@@
Temporal Correspondence: Distinct frames in dynamic point cloud sequences exhibit varying geometries—different noise patterns, point counts, and no guaranteed point-to-point correspondence. FlowNet3D [LQG19] pioneered end-to-end scene flow learning, estimating vector displacements between consecutive frames without assuming point-to-point correspondence. St4RTrack [FZW25] creates dynamic point cloud sequence from videos and tracks point motion through an attention-enabled tracking branch supported by video inputs. Alternatively, correspondence can be enforced through interpolation between adjacent frames. FastPCI [ZQXY24] applies pyramidal transformer blocks to extract motion and structure features facilitating in-betweening generation.
Appearance Modeling: The discrete, isolated nature of point clouds prevents complete appearance modeling, leaving void spaces between points unlike surface or volumetric representations. PAPR-in-Motion [PZL24] addresses this by associating each point with view-dependent features and rendering through attention-weighted feature interpolation, achieving void-free rendering while enabling dynamic visualization through geometric point displacement.
Open Challenges. Point clouds remain the primary interface with real-world sensor captures (LiDAR, RGB-D cameras), and this abundance of point-based training data has recently enabled pointmap-based methods to achieve impressive 4D reconstruction from monocular videos. However, a reconstruction-generation asymmetry persists: while point clouds excel at capturing dynamic geometry from observations, their inherent limitations—discrete sampling without explicit surface connectivity—make them less suitable for generation tasks. Incorporating physical motion remains challenging: the lack of topological structure complicates enforcement of physical constraints such as collision response, conservation laws, and contact dynamics. Future work should focus on efficient bidirectional conversion between point clouds and other representations, enabling knowledge transfer of mesh topology, NeRF appearance modeling, and physical constraints to point-based frameworks while preserving their computational efficiency and sensor correspondence advantages.
@@ 2.1.3 Neural Radiance Fields (NeRF) based Representation
Neural Radiance Fields (NeRF) [MST21] represent a 3D scene by mapping a 3D query point and view direction to density and color using an MLP. The final pixel color is obtained by integrating these color predictions along camera rays via volume rendering [KVH84]. This continuous representation achieves high visual fidelity from only 2D images and can capture topologically complex phenomena such as smoke and fluids, and storing dynamic scenes compactly within an MLP [PSH21, AHR23]. Despite its promise for modeling dynamic 3D scenes, NeRF faces several key challenges.
@@
Flickering and Unrealistic Deformations: A direct extension of NeRF to the 4D domain models temporal variations by conditioning the representation on a time signal or a learnable embedding for each timestamp in an auto-decoder manner [GSKH21, LNSW21, LSZ22, WZT22, OMT21, SCL23]. Each frame’s color and density are then composited and rendered using the standard NeRF volume rendering framework. However, this approach often suffers from temporal flickering and unrealistic deformations. These are typically alleviated through temporal consistency losses and rigidity constraints [GSKH21, LNSW21].
@@
Slow Convergence: NeRF-based representations often suffer from slow convergence due to the heavy computational cost of training large MLPs for dynamic scenes. To accelerate training, several methods introduce learnable spatial-temporal data structures such as Hex-planes and K-planes [CJ23, SSP23, FKMW23, JZG24], which trade memory for speed. These approaches sample positional features from planar decompositions via bilinear interpolation before passing them through a smaller MLP for color and density prediction. By offloading computation from the MLP to explicit plane-based storage, convergence accelerates significantly compared to standard NeRFs. However, planar decompositions may sacrifice fine-grained spatial details. Alternative data structures have also been explored, including voxel grids [LCM22, FYW22, SZT23, YYPZ23] that provide denser spatial sampling, and hash grids [BSR24, ZYX23, ZCW24] that offer adaptive resolution through learned hash encodings.
@@
Handling Sparse Input Scenarios: NeRF representation, in general, requires dense multi-view video captures as input, limiting their applicability. To improve robustness under sparse input conditions, recent works incorporate flow supervision [WMJL23, YHR24] to train models with only monocular video supervision. Other methods approach this problem through motion-adjusted feature aggregation [LWC23] or enriching input data from generative models [YXV25, XYV24].
@@
Open Challenges. Despite these improvements, temporal flickering and unrealistic motion artifacts remain persistent issues, particularly when training from sparse or limited input views. Motion synthesis and editing in NeRF-based methods are also less intuitive, as manipulations occur in latent deformation spaces rather than through direct geometric controls. Additionally, NeRF’s ray marching mechanism incurs significant computational overhead during inference, leading the community to increasingly favor Gaussian Splatting ( Section˜2.1.4) as a more efficient alternative for real-time dynamic scene rendering.

### v2

@@ Abstract
We present a survey on 4D generation and reconstruction, a fast-evolving subfield of computer graphics whose developments have been propelled by recent advances in neural fields, geometric and motion deep learning, as well as 3D generative artificial intelligence (GenAI). While our survey is not the first of its kind, we build our coverage of the domain from a unique and distinctive perspective of 4D representations, to model 3D geometry evolving over time while exhibiting motion and interaction. Specifically, instead of offering an exhaustive enumeration of many works, we take a more selective approach by focusing on representative works to highlight both the desirable properties and ensuing challenges of each representation under different computation, application, and data scenarios. The main take-away message we aim to convey to the readers is on how to select and then customize the appropriate 4D representations for their tasks. Organizationally, we separate the 4D representations based on three key pillars: geometry, motion, and interaction. Our discourse will not only encompass the most popular representations of today, such as neural radiance fields (NeRFs) and 3D Gaussian Splatting (3DGS), but also bring attention to relatively under-explored representations in the 4D context, such as structured models and long-range motions. Throughout our survey, we will reprise the role of large language models (LLMs) and video foundational models (VFMs) in a variety of 4D applications, while steering our discussion towards their current limitations and how they can be addressed. We also provide a dedicated coverage on what 4D datasets are currently available, as well as what is lacking, in driving the subfield forward.
@@ 1 Introduction
Recent surveys in this domain have primarily categorized methods by applications, approaches, and/or the granularity of the extracted scene signals. Cao et al. [CLH25] provides a structured review of recent progress in reconstructing 4D spatial intelligence, which is defined as understanding 3D scenes and their evolution over time, advancing from extracting low-level geometric cues such as depth, pose, and point maps toward complex reasoning encompassing interactions and physics. This survey does highlight the growing importance of advanced 3D representations while focusing mainly on neural radiance fields (NeRFs) and 3D Gaussian splatting (3DGS). Along the same lines, Fan et al. [FZZ25] presents a comprehensive survey of recent progress in dynamic scene representation and reconstruction, focusing on the transition from NeRF to 3DGS and their trade-offs. This second survey systematically categorizes the works covered based on motion representation paradigms including rigid, articulated, non-rigid, and hybrid motions, and examines strategies for modeling temporal changes in geometry and appearance. In addition, Miao et al. [MLQ25] survey the emerging field of 4D generation by introducing a taxonomy of low-level geometry representations (meshes, NeRFs, point clouds, and 3DGS), foundational techniques (diffusion models and score distillation sampling), and conditioning methods (text, images, videos, 3D inputs, and multimodal control). This survey further categorizes the algorithmic approaches (end-to-end, via generated data, implicit distillation, vs. explicit supervision) and highlights the growing range of applications for 4D generation.
@@
Organizationally, we separate the 4D representations based on three key pillars: geometry, motion, and interaction. We distinguish between structured representations, i.e., those that maintain explicit primitive or part delineations and relations, and unstructured representations, i.e., those that encode 4D content through implicit functions, point distributions, or learned feature spaces without explicit structural constraints. While the latter category encompasses the most popular representations of today, namely NeRFs and 3DGS, they are both built on rendering primitives, mainly for novel view synthesis. They are not best suited to prevalent 4D tasks such as modeling, editing, or interactions, for which more compact and structural representations, built on higher-level primitives, are more appropriate. Hence, we also discuss these representational choices even though they have been relatively under-explored. Throughout our survey, we will reprise the role of large language models (LLMs) and video foundational models (VFMs) in a variety of 4D applications, while steering our discussion towards their current limitations and how they can be addressed. We also provide a dedicated coverage on what 4D datasets are currently available, as well as what is lacking, in driving the subfield forward.
@@
Our survey is organized as follows. Section˜2 presents our core representational analysis, examining unstructured representations (mesh, point clouds, NeRFs and 3DGS) and structured representations (template, parts, and graphs). Section˜3 focuses on motion and temporal dynamics, analyzing how different motion types interact with representational choices. Section˜4 covers representation choices in modeling multiple entities that interact with each other. Section˜5 surveys datasets, evaluation metrics, and benchmarking frameworks that enable systematic comparison across representations. Section˜6 provides an overview on training paradigms. Finally, Sections˜7 and 8 provide an overall analysis across representations and discuss emerging trends, persistent challenges, and future research directions in 4D representation. The overall taxonomy of our survey is shown in Figure˜2.
@@ 2 Modeling Geometry
We categorize representations into unstructured (Section˜2.1) and structured (Section˜2.2) classes based on whether their operational primitives carry explicit functional, hierarchical, or semantic meaning. See Figure˜1 for an illustration. Unstructured representations—meshes (Section˜2.1.1), point clouds (Section˜2.1.2), NeRF (Section˜2.1.3), and Gaussian Splatting (Section˜2.1.4)—use primitives (vertices, points, ray samples, Gaussians) as independent geometric elements. While local connectivity may exist (e.g., mesh faces), these representations do not impose inter-primitive global structural constraints or functional decompositions.
@@ 2.1.1 Mesh-Based Representations
Meshes are mainly adopted in object-centric 4D modeling, aimed at producing animated mesh assets from text [DSW25, CHC24], video [CZTW25, LCL24], or static 3D assets [WYWB25, SLW25]. However, mesh representations face several challenges in 4D applications, which we discuss below alongside methods developed to address them.
Spatiotemporal Consistency: Animated mesh sequences require precise vertex-to-vertex correspondence across frames to preserve consistent texture and topology, preventing naive per-frame generation pipelines that can lead to different vertex-connectivity in each generated sample. V2M4 [CZTW25] circumvents this through a reconstruction-and-refine pipeline: it first applies image-to-3D reconstruction models [XLX25, ZLL25, LZL25, LLY24] to generate per-frame meshes, then sets the first frame as an anchor and deforms it to register with subsequent frames, maintaining consistent mesh correspondence through continuous deformation.
@@
Open Challenges: Despite recent progress, fundamental limitations remain. Topological inflexibility prevents meshes from representing splitting or merging. Their inflexible topology and connectivity struggle with volumetric phenomena such as clouds, smoke, and fire, motivating the adoption of topologically flexible alternatives such as NeRF (Section˜2.1.3) and point clouds (Section˜2.1.2) for such scenarios. Additionally, current mesh-centric generation and reconstruction methods predominantly focus on object-level rather than scene-level modeling. The latter has recently been studied on static 3D scene reconstruction [YZY25] through generative methods, while most existing works rely on mesh database retrieval [WIR24]. Key difficulties include: (1) generating and tracking separate mesh topologies for multiple objects at varying scales, (2) handling inter-mesh physical constraints, and (3) maintaining temporal coherence when objects move independently with different deformation patterns. Addressing these challenges remains an open problem for the research community.
@@ 2.1.2 Point-cloud-based Representation
A point cloud represents 3D geometry as an unordered set of points $P=\{p_{i}\}_{i=1}^{N}$, where each point $p_{i}$ is defined by spatial coordinates $(x,y,z)$ and optional attributes including color (RGB), intensity, surface normals, and timestamps. Their unstructured, unordered nature enables representation of arbitrary and changing topologies without connectivity constraints. Point clouds also serve as the primary output format for real-world acquisition mechanisms. This direct sensor correspondence has established point clouds as the standard representation for large-scale scene capture, particularly in autonomous driving (KITTI [GLSU13], nuScenes [CBL20]), where their geometric scalability outperforms alternatives. On native 4D point cloud sequences, numerous works focus on improving point cloud quality through denoising [HH25], upsampling [LLF21, BBBB23], and mesh reconstruction [TADTBS22, RBZ20, NMOG19]. Another research direction addresses dynamic point cloud understanding, including motion recognition [LHL25, DZLY23], object detection and tracking [HGH22, WNGO24, CLZ24], scene flow estimation [LQG19, WWL20, TD20], forecasting [KHHR23, WYW24], and motion interpolation [ZQZ22, ZWL23, LCQ21]. With the emergence of large feed-forward geometric models like DUST3R [WLC24] and VGGT [WCK25], a new promising branch of work is to create dynamic point map sequences from videos [ZHH24, WZH25, WZZ25, TZW25] and incorporate tracking within the inferred point maps [XWX25].
@@
Temporal Correspondence: Different frames in dynamic point cloud sequences exhibit varying noise patterns, point counts, and no guaranteed point-to-point correspondence. FlowNet3D [LQG19] pioneered end-to-end scene flow learning, estimating vector displacements between consecutive frames without assuming point-to-point correspondence. St4RTrack [FZW25] generates dynamic point cloud sequences from videos and tracks point motion using an attention-enabled tracking branch that leverages video inputs. Alternatively, correspondence can be enforced by interpolating between adjacent frames. FastPCI [ZQXY24] applies pyramidal transformer blocks to extract motion and structure features, facilitating in-betweening generation.
Appearance Modeling: The discrete, isolated nature of point clouds prevents complete appearance modeling, leaving void spaces between points unlike surface or volumetric representations. PAPR-in-Motion [PZL24] addresses this by associating each point with view-dependent features and rendering through attention-weighted feature interpolation, achieving void-free rendering while enabling dynamic visualization through point displacement.
Open Challenges: Point clouds remain the primary interface with real-world sensor captures (LiDAR, RGB-D cameras), and this abundance of point-based training data has recently enabled pointmap-based methods to achieve impressive 4D reconstruction from monocular videos. However, a reconstruction-generation asymmetry persists: while point clouds excel at capturing dynamic geometry from observations, their inherent limitations—discrete sampling without explicit surface connectivity—make them less suitable for generation tasks. Incorporating physical motion remains challenging: the lack of topological structure complicates enforcement of physical constraints such as collision response, conservation laws, and contact dynamics. Future work should focus on efficient bidirectional conversion between point clouds and other representations, enabling knowledge transfer of mesh topology, NeRF appearance modeling, and physical constraints to point-based frameworks while preserving their computational efficiency and sensor correspondence advantages.
@@ 2.1.3 Neural Radiance Fields (NeRF) based Representation
Neural Radiance Fields (NeRF) [MST20] represent 3D scenes by mapping 3D query points and view directions to density and color using an MLP. Final pixel colors are obtained by integrating these color predictions along camera rays via volume rendering [KVH84]. This continuous representation achieves high visual fidelity from only 2D images and can capture topologically complex phenomena like smoke and fluids, and compactly store dynamic scenes within an MLP [PSH21, AHR23]. Despite its promise, NeRF faces several key challenges in 4D modeling.
@@
Flickering and Unrealistic Deformations: A direct extension of NeRF to the 4D domain models temporal variations by introducing time-varying fields. This is typically done by conditioning the representation on a time signal or by learning a unique embedding for each timestamp  [GSKH21, LSZ22, WZT22, OMT21]. Another approach models dynamics by predicting deformation or flow fields using a separate deformation network, which warps the scene over time  [LNSW21, SCL23]. Each frame’s color and density are then composited and rendered using the standard NeRF volume rendering framework. However, this approach often suffers from temporal flickering and unrealistic deformations. These issues are typically alleviated through temporal consistency losses and rigidity constraints [GSKH21, LNSW21].
@@
Slow Convergence: NeRF-based representations often suffer from slow convergence due to the heavy computational cost of training large MLPs for dynamic scenes. To accelerate training, several methods introduce learnable spatial-temporal data structures such as Hex-planes and K-planes [CJ23, SSP23, FKMW23, JZG24], which trade memory for speed. These approaches sample positional features from planar decompositions via bilinear interpolation before passing them through a smaller MLP for color and density prediction. By offloading computation from the MLP to explicit plane-based storage, convergence accelerates significantly compared to standard NeRFs. However, planar decompositions may sacrifice fine-grained spatial details. Alternative data structures have also been explored, including voxel grids [LCM22, FYW22, SZT23, YYPZ24] that provide denser spatial sampling, and hash grids [BSR24, ZYX23, ZCW24] that offer adaptive resolution through learned hash encodings.
@@
Handling Sparse Input Scenarios: NeRF representations, in general, require dense multi-view video captures as input. To improve robustness under sparse input conditions, recent works incorporate flow supervision [WMJL23, YHR24] to train models with only monocular video supervision. Other methods approach this problem through motion-adjusted feature aggregation [LWC23] or enriching input data from generative models [YXV25, XYV24].
@@
Open Challenges: Despite these improvements, temporal flickering and unrealistic motion artifacts remain persistent issues, particularly when training from sparse or limited input views. Motion synthesis and editing in NeRF-based methods are also less intuitive, as manipulations occur in latent deformation spaces rather than through direct geometric controls. Additionally, NeRF’s ray marching mechanism incurs significant computational overhead during inference, leading the community to increasingly favor Gaussian Splatting(Section˜2.1.4) as a more efficient alternative for real-time dynamic scene rendering.